Two Computational Challenges About Chomp Shalosh B. EKHAD and Doron ZEILBERGER

Dedicated to Richard K. Guy (b. Sept. 30, 1916) in honor of his forthcoming

102th birthday

0 1 × 1 1 × 2 1 × 3 1 × 4 1 × 5 1 × 6 1 × 7 1 × 8 1 × 9 1 × 10 1 × 11 1 × 12 1 × 13